For a body moving along a straight line, the displacement is given by the equation s= a +bt + ct^2 + d^t3 where c=3ms^2 and d=2ms^3 Find the time when acceleration becomes 30ms^2

Grade:11

1 Answers

Khimraj
5 years ago
s = a + bt + ct2 + dt3
v = ds/dt = b + 2ct + 3dt2
a = dv/dt = 2c + 6dt
c = 3,d =2, acceleration = 30
30 = 2*3 + 6*2*t
24 = 12t
So t = 2 sec
Hope it clears. If you like anser then please approve it.